a+2b+0c-3d=0,2a-b+c+7d=0,4a+3b+2c+2d=0 solve this by cramers rule
Answer by richwmiller(17219)

a+2b+0c-3d=0
2a-b+c+7d=0
4a+3b+2c+2d=0
1,2,0,-3,0
2,-1,1,7,0
4,3,2,2,0
4 unknowns with only three equations.
Sorry, can't be solved by Cramer's rule.
b = -4a/3, c = 5a/9, d =-5a/9
a = 9n, b = -12n, c = 5n, d = -5n
possible solutions
a = 9 , b = -12 , c = 5 , d = -5
a = 18, b = -24, c = 10, d = -10